हिंदी

The probability that the home team will win an upcoming football game is 0.77, the probability that it will tie the game is 0.08, and the probability that it will lose the game is ______.

Advertisements
Advertisements

प्रश्न

The probability that the home team will win an upcoming football game is 0.77, the probability that it will tie the game is 0.08, and the probability that it will lose the game is ______.

रिक्त स्थान भरें
Advertisements

उत्तर

The probability that the home team will win an upcoming football game is 0.77, the probability that it will tie the game is 0.08, and the probability that it will lose the game is 0.15.

Explanation:

P(Loosing the game) = 1 – (0.77 + 0.08)

= 1 – 0.85

= 0.15

shaalaa.com
  क्या इस प्रश्न या उत्तर में कोई त्रुटि है?
अध्याय 16: Probability - Exercise [पृष्ठ ३०१]

APPEARS IN

एनसीईआरटी एक्झांप्लर Mathematics Exemplar [English] Class 11
अध्याय 16 Probability
Exercise | Q 37 | पृष्ठ ३०१

वीडियो ट्यूटोरियलVIEW ALL [1]

संबंधित प्रश्न

A coin is tossed twice, what is the probability that at least one tail occurs?


A card is selected from a pack of 52 cards.

  1. How many points are there in the sample space?
  2. Calculate the probability that the card is an ace of spades.
  3. Calculate the probability that the card is
    1. an ace
    2. black card.

A fair coin with 1 marked on one face and 6 on the other and a fair die are both tossed. Find the probability that the sum of numbers that turn up is (i) 3 (ii) 12


A fair coin is tossed four times, and a person win Re 1 for each head and lose Rs 1.50 for each tail that turns up.

From the sample space calculate how many different amounts of money you can have after four tosses and the probability of having each of these amounts.


Three coins are tossed once. Find the probability of getting

  1. 3 heads
  2. 2 heads
  3. at least 2 heads
  4. at most 2 heads
  5. no head
  6. 3 tails
  7. exactly two tails
  8. no tail
  9. atmost two tails.

Fill in the blank in following table:

P(A) P(B) P(A ∩ B) P(A ∪ B)
`1/3` `1/5` `1/15` ....

4 cards are drawn from a well-shuffled deck of 52 cards. What is the probability of obtaining 3 diamonds and one spade?


If 4-digit numbers greater than 5,000 are randomly formed from the digits 0, 1, 3, 5, and 7, what is the probability of forming a number divisible by 5 when, the digits are repeated?


A and B throw a pair of dice. If A throws 9, find B's chance of throwing a higher number.

 

Two unbiased dice are thrown. Find the probability that  neither a doublet nor a total of 8 will appear


A bag contains 8 red, 3 white and 9 blue balls. If three balls are drawn at random, determine the probability that all the three balls are blue balls 


Which of the cannot be valid assignment of probability for elementary events or outcomes of sample space S =  {w1, w2, w3, w4, w5, w6, w7}:

Elementary events: w1 w2 w3 w4 w5 w6 w7
(ii)
\[\frac{1}{7}\]
\[\frac{1}{7}\]
\[\frac{1}{7}\]
\[\frac{1}{7}\]
\[\frac{1}{7}\]
\[\frac{1}{7}\]
\[\frac{1}{7}\]

Which of the cannot be valid assignment of probability for elementary events or outcomes of sample space S =  {w1, w2, w3, w4, w5, w6, w7}:

Elementary events: w1 w2 w3 w4 w5 w6 w7
(iv)
\[\frac{1}{14}\]
\[\frac{2}{14}\]
\[\frac{3}{14}\]
\[\frac{4}{14}\]
\[\frac{5}{14}\]
\[\frac{6}{14}\]
\[\frac{15}{14}\]

A box contains 100 bulbs, 20 of which are defective. 10 bulbs are selected for inspection. Find the probability that: all 10 are defective


A box contains 100 bulbs, 20 of which are defective. 10 bulbs are selected for inspection. Find the probability thatat least one is defective


A box contains 100 bulbs, 20 of which are defective. 10 bulbs are selected for inspection. Find the probability that none is defective


Two dice are thrown together. The probability that neither they show equal digits nor the sum of their digits is 9 will be


Three squares of chessboard are selected at random. The probability of getting 2 squares of one colour and other of a different colour is ______.


Six new employees, two of whom are married to each other, are to be assigned six desks that are lined up in a row. If the assignment of employees to desks is made randomly, what is the probability that the married couple will have nonadjacent desks?


The accompanying Venn diagram shows three events, A, B, and C, and also the probabilities of the various intersections (for instance, P(A ∩ B) = .07). Determine P(A)


The accompanying Venn diagram shows three events, A, B, and C, and also the probabilities of the various intersections (for instance, P(A ∩ B) = .07). Determine `P(B ∩ barC)`


The accompanying Venn diagram shows three events, A, B, and C, and also the probabilities of the various intersections (for instance, P(A ∩ B) = .07). Determine `P(A ∩ barB)`


The accompanying Venn diagram shows three events, A, B, and C, and also the probabilities of the various intersections (for instance, P(A ∩ B) = .07). Determine P(B ∩ C)


The accompanying Venn diagram shows three events, A, B, and C, and also the probabilities of the various intersections (for instance, P(A ∩ B) = .07). Determine Probability of exactly one of the three occurs


A bag contains 8 red and 5 white balls. Three balls are drawn at random. Find the probability that all the three balls are red


A bag contains 8 red and 5 white balls. Three balls are drawn at random. Find the probability that all one ball is red and two balls are white


If the letters of the word ASSASSINATION are arranged at random. Find the probability that four S’s come consecutively in the word


In a non-leap year, the probability of having 53 tuesdays or 53 wednesdays is ______.


While shuffling a pack of 52 playing cards, 2 are accidentally dropped. Find the probability that the missing cards to be of different colours ______.


The probability that a person visiting a zoo will see the giraffee is 0.72, the probability that he will see the bears is 0.84 and the probability that he will see both is 0.52.


The probability that a student will pass his examination is 0.73, the probability of the student getting a compartment is 0.13, and the probability that the student will either pass or get compartment is 0.96.


If A and B are two candidates seeking admission in an engineering College. The probability that A is selected is .5 and the probability that both A and B are selected is at most .3. Is it possible that the probability of B getting selected is 0.7?


Share
Notifications

Englishहिंदीमराठी


      Forgot password?
Use app×